OpenStudy (anonymous):

soybean meal is 18% protein cornmeal is 9% protein how many pound of each should be mixed together in order to get 360lb mixture that is 16% protein

OpenStudy (anonymous):

god i love this problem. put x = number of pounds of soybean. since the total is 360 then the amount of cornmeal must be \[360-x\] since they have to total 360

OpenStudy (anonymous):

you get 18% protein from soybean, so it contributes \[.18x\] and likewise the cornmeal contributes \[.09(360-x)\]

OpenStudy (anonymous):

when you combine them you have to get 16% of 360 = \[.16\times 360=57.6\] so set \[.18x+.09(360-x)=57.6\] and solve for x

OpenStudy (anonymous):

so it would be x=57.6

OpenStudy (anonymous):

oh no, we still have to solve \[.18x+.09(360-x)=57.6\] for x

OpenStudy (anonymous):

i would multiply everything by 100 to get rid of the annoying decimals and write \[18x+9(360-x)=5760\]

OpenStudy (anonymous):

then \[18x+3240-9x=5760\] \[9x+3240=5760\] \[9x=2520\] and finally \[x=\frac{5250}{9}=580\]

OpenStudy (anonymous):

hold the phone i made a mistake let me find it

OpenStudy (anonymous):

oh right last line should be \[x=\frac{2520}{9}=280\]

OpenStudy (anonymous):

so 280 pounds of soy, the rest cornmeal

OpenStudy (anonymous):

cornmeal would be the same